OpenStudy (anonymous):

Solve the equation. (If there is no solution, enter NO SOLUTION.) 4 x − 3 = 3 x + 5

OpenStudy (godlovesme):

4x-3=3x+5 is that the eqn? @simonej46

OpenStudy (anonymous):

no it's a fraction

OpenStudy (anonymous):

well yes

OpenStudy (lizz123):

\[\frac{ 4 }{ x }-3=\frac{ 3 }{ x }+5\]

OpenStudy (godlovesme):

like @lizz123 's ? @simonej46

OpenStudy (anonymous):

the 4 is on top x-3 on the bottom a = 3 on top and x+5 on the bottom

OpenStudy (lizz123):

so is it how i wrote it

OpenStudy (anonymous):

yes

OpenStudy (lizz123):

\[x =\frac{ 1 }{ 8 }\] I dont know if I am correct

OpenStudy (anonymous):

k

OpenStudy (lizz123):

k